OpenStudy (sasogeek):

please solve in radians from start to finish, even if u won't explain, please show working. 7sec(3t)=-10

OpenStudy (nottim):

Divide both by 7.

OpenStudy (nottim):

Then, do what sec does. (Reverse of whatever is it). This should the right side -1/70 after 2 steps.

OpenStudy (nottim):

After, divide both sides by 3

OpenStudy (nottim):

I was incredibly vague in step 2. IF you'd like me to refine that, tell me. But I am busy.

OpenStudy (nottim):

Is sec 1/cos?

OpenStudy (sasogeek):

yes

OpenStudy (nottim):

1. Divide both sides by 7. 2.sec3t=-10/7 3.1/cos3t=-10/7 4. cos3t=7/-10 5.cos3t/3=(7/-10)/3 6.t=cos^-1[(7/-10)/3 DERONE!

OpenStudy (nottim):

HOOAH

OpenStudy (mertsj):

sec(3t)=-10/7 cos(3t)=-7/10 3t = 2.346 +2pi=8.629 +2pi=14.912 t=.782, 2.876,4.971 3t = 3.937 +2pi=10.220+2pi=16.503 t=1.312,3.407,5.501

OpenStudy (nottim):

AWw. Sigh.

OpenStudy (mertsj):

5.cos3t/3=(7/-10)/3 YOU CAN'T DO THAT!!!!

OpenStudy (sasogeek):

Mertsj's answers are the same as that in my book... i guess i was on the right path... i got stuck after getting the angle 3t.... since that was the angle, i wasn't sure if i had to continue to find t or they just needed the angle. thanks a ton :)

OpenStudy (nottim):

That's why I failed that test.

OpenStudy (nottim):

Eroops.